OpenStudy (anonymous):

7*7^2/7^9 use properties of exponents to simplify

OpenStudy (mertsj):

\[\frac{7(7^2)}{7^9}=\frac{7^3}{7^9}=\frac{1}{7^6}=\frac{1}{117649}\]